A machine’s motion is as shown below. NOTE: The line is NOT straight between t=6 and t=8s.

1) Draw the s-t graph and a-t graph. Label key points, axes; clearly distinguish between straight and curved lines. Enter the value of acceleration, a, at t=8 s, for the online answer below.

2) Determine the function of acceleration (a) and distance (s) in terms of time (t) over the time interval 6 to 8 seconds.

18 15 12 9 6 velocity (ft/s) v = 36-0.52 (from t=6 to t=8) 0 -6 -9 a time (5)
Answer #1

Solution is velocity-time graph is given ve rate of change of position (displacement) with respect to time . V- 52-51 ds at ttu 2 to 5 sec. V u a 9 straight line yayı ine ma 18-(-9) 5-2 m (x-2) :: V- (9):9 (t-2) V +92 94 -18 vaqt-27 m/s -) t = 5 to 6UST we know that Vads di- dsavat Sasa fudt 5= fvat te Where ca constant of integratin. t=0 to 2 sec S so fyat te=f-4.5 t) dtPutting in SP 9t² 27 t tc 2. 2) -9 = 9(2) 9 -27*2 + c2 5) C 27 saqt² -27t + 27 sö 4.5+? - 27t +27 for to 2 to 5 sec s(t-5) = 4.5*25-27 x5+27 2 4.5 m t= sto osec so sudt to Sisat5436) - 22.5m 57 22.5 = 36 x 6 -0.5 X 6 ac ६ Sit Sleep :: Ś= 367 + С - ) S 6 0.5t 3 36t t3 - 157.5 G scat t=85). 36x8-83 -157Sul 45.166 such 22.5 5 (m) 14:30 2 5 3 6 ܝܢ 7 8 tes -१a av dt dur for t=0 to 25 ae dl-4.5+) -4.5 m/s² at a 2 -4.5 ups 2 for to t² n 2 tussec १ aad (qt-27) at az 9-02 aaq mys2 forQ2 aso-0.5x2t et -t aart! at ²6 as-buf? UY (es/ne) 0 > t (sec) J 5 -4.5 -6 -8 (2) for t-6 to t-8 Ta=-t] (already found before